OpenStudy (anonymous):

A company has plans to increase its sales at the rate of 8% per year. If the sales of the company in 2001 were $5 million, how much sales can you expect the company to make in the year 2003?

OpenStudy (anonymous):

2003-2001=2 years A=5(1+0.08)^2=6.29856 round: 6.3 Million

OpenStudy (anonymous):

thank you for the medal

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

but i got a different answer i got 5830000

OpenStudy (anonymous):

Here is the equation. Hope it helps\[5 \times 10^{6} \times 1.08^{n}\] where n is the number of years the business has been running. Bonus help : in 2001, n = 0. There you go :)

OpenStudy (anonymous):

I cacul yes ,you r correct sorry 5.8Million